# Fact-or-Fair: A Checklist for Behavioral Testing of AI Models on Fairness-Related Queries

### v1

@@ Acknowledgments
We would like to thank Professor Jieyu Zhao from University of Southern California for her valuable suggestions during this research. The work is supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK14206921 of the General Research Fund).
Appendix A Proof of the Accuracy-Entropy Trade-Off
@@
$-\sum_{\begin{subarray}{c}j=1,\cdots,k\\ j\neq i\end{subarray}}p_{j}\log p_{j}-a\log a,$ (1)
@@
$\mathcal{L}(p_{1},\dots,p_{i-1},p_{i+1},\dots,p_{k},\lambda)=-\sum_{\begin{subarray}{c}j=1,\cdots,k\\ j\neq i\end{subarray}}p_{j}\log p_{j}+\lambda\left(\sum_{\begin{subarray}{c}j=1,\cdots,k\\ j\neq i\end{subarray}}p_{j}-(1-a)\right).$ (3)
@@ Appendix B Quantitative Results
Gender GPT-3.5-Turbo-0125 84.44 53.33 67.24 53.17 53.35 Midjourney 48.90 51.10
@@
Race GPT-3.5-Turbo-0125 39.81 33.33 48.78 28.71 30.73 Midjourney 25.36 22.36
@@
Gender GPT-3.5-Turbo-0125 21.43 99.86 94.10 99.98 99.96 Midjourney 96.25 99.00
@@
Race GPT-3.5-Turbo-0125 13.49 97.80 90.34 99.16 97.80 Midjourney 81.65 75.99
@@
Gender GPT-3.5-Turbo-0125 21.43 97.45 83.88 98.88 98.58 Midjourney 64.36 74.43
@@
Race GPT-3.5-Turbo-0125 13.49 92.96 83.12 95.71 93.02 Midjourney 55.53 55.32
@@
(a) LLM O S-B S-R S-A S-G (b) T2I Model O S
Gender GPT-3.5-Turbo-0125 $<10^{-6}$ 94.66 63.4 97.79 96.99 Midjourney 89.48 96.10
GPT-4o-2024-08-06 $<10^{-6}$ 93.54 64.28 93.82 91.04 DALL-E 3 57.98 71.26
Gemini-1.5-Pro $<10^{-6}$ 94.75 60.31 93.95 94.78 SDXL-Turbo 88.33 91.91
LLaMA-3.2-90B-Vision-Instruct $<10^{-6}$ 96.22 65.77 97.49 97.25 Flux-1.1-Pro 91.33 74.64
WizardLM-2-8x22B $<10^{-6}$ 95.82 73.26 96.13 95.30
Qwen-2.5-72B-Instruct $<10^{-6}$ 95.65 67.62 96.85 96.33
Race GPT-3.5-Turbo-0125 $<10^{-6}$ 68.77 42.76 80.50 68.52 Midjourney 58.73 46.26
GPT-4o-2024-08-06 $<10^{-6}$ 75.34 39.75 75.18 71.43 DALL-E 3 17.67 40.12
Gemini-1.5-Pro $<10^{-6}$ 77.42 58.43 77.92 73.74 SDXL-Turbo 31.23 57.52
LLaMA-3.2-90B-Vision-Instruct $<10^{-6}$ 75.83 51.56 80.06 73.51 Flux-1.1-Pro 39.82 30.29
WizardLM-2-8x22B $<10^{-6}$ 73.51 53.00 81.48 72.39
Qwen-2.5-72B-Instruct $<10^{-6}$ 75.12 41.61 82.92 71.11
@@
Gender GPT-3.5-Turbo-0125 11.89 2.18 4.80 0.82 1.07 4.15 Midjourney 29.14 23.27 26.21
@@
Race GPT-3.5-Turbo-0125 53.17 5.51 5.79 3.99 6.21 14.93 Midjourney 41.97 44.05 43.01
@@ Appendix E Racial Information in the Statistics
Economic Employment Rate Female, Male Asian, Black, Hispanic, White
@@
Social Educational Attainment Female, Male Asian, Black, Hispanic, White
@@
Health Life Expectancy Female, Male Asian, Black, Hispanic, White

### v3

@@ Acknowledgments
We would like to thank Professor Jieyu Zhao from University of Southern California for her valuable suggestions for this research. The work is supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK14206921 of the General Research Fund).
Appendix A Proof of the Upper-Bound
@@
$\sum_{\begin{subarray}{c}j=1,\cdots,k\\ j\neq i\end{subarray}}p_{j}\log p_{j}-a\log a,$ (1)
@@
$\mathcal{L}(p_{1},\dots,p_{i-1},p_{i+1},\dots,p_{k},\lambda)=$ (3)
$-\sum_{\begin{subarray}{c}j=1,\cdots,k\\ j\neq i\end{subarray}}p_{j}\log p_{j}+\lambda\left(\sum_{\begin{subarray}{c}j=1,\cdots,k\\ j\neq i\end{subarray}}p_{j}-(1-a)\right).$ (4)
@@ Appendix B Quantitative Results
Gender
GPT-3.5-Turbo-0125 84.44 53.33 67.24 53.17 53.35 Midjourney 48.90 51.10
@@
Race
GPT-3.5-Turbo-0125 39.81 33.33 48.78 28.71 30.73 Midjourney 25.36 22.36
@@
Gender
GPT-3.5-Turbo-0125 21.43 99.86 94.10 99.98 99.96 Midjourney 96.25 99.00
@@
Race
GPT-3.5-Turbo-0125 13.49 97.80 90.34 99.16 97.80 Midjourney 81.65 75.99
@@
Gender
GPT-3.5-Turbo-0125 21.43 97.45 83.88 98.88 98.58 Midjourney 64.36 74.43
@@
Race
GPT-3.5-Turbo-0125 13.49 92.96 83.12 95.71 93.02 Midjourney 55.53 55.32
@@
(a) LLM O S-B S-R S-A S-G (b) T2I Model O S
Gender
GPT-3.5-Turbo-0125 $10^{-6}$ 94.66 63.4 97.79 96.99 Midjourney 89.48 96.10
GPT-4o-2024-08-06 $10^{-6}$ 93.54 64.28 93.82 91.04 DALL-E 3 57.98 71.26
Gemini-1.5-Pro $10^{-6}$ 94.75 60.31 93.95 94.78 SDXL-Turbo 88.33 91.91
LLaMA-3.2-90B-Vision-Instruct $10^{-6}$ 96.22 65.77 97.49 97.25 Flux-1.1-Pro 91.33 74.64
WizardLM-2-8x22B $10^{-6}$ 95.82 73.26 96.13 95.30
Qwen-2.5-72B-Instruct $10^{-6}$ 95.65 67.62 96.85 96.33
Race
GPT-3.5-Turbo-0125 $10^{-6}$ 68.77 42.76 80.50 68.52 Midjourney 58.73 46.26
GPT-4o-2024-08-06 $10^{-6}$ 75.34 39.75 75.18 71.43 DALL-E 3 17.67 40.12
Gemini-1.5-Pro $10^{-6}$ 77.42 58.43 77.92 73.74 SDXL-Turbo 31.23 57.52
LLaMA-3.2-90B-Vision-Instruct $10^{-6}$ 75.83 51.56 80.06 73.51 Flux-1.1-Pro 39.82 30.29
WizardLM-2-8x22B $10^{-6}$ 73.51 53.00 81.48 72.39
Qwen-2.5-72B-Instruct $10^{-6}$ 75.12 41.61 82.92 71.11
@@
Gender
GPT-3.5-Turbo-0125 11.89 2.18 4.80 0.82 1.07 4.15 Midjourney 29.14 23.27 26.21
@@
Race
GPT-3.5-Turbo-0125 53.17 5.51 5.79 3.99 6.21 14.93 Midjourney 41.97 44.05 43.01
@@ Appendix E Racial Information in the Statistics
Economic
Employment Rate Female, Male Asian, Black, Hispanic, White
@@
Social
Educational Attainment Female, Male Asian, Black, Hispanic, White
@@
Health
Life Expectancy Female, Male Asian, Black, Hispanic, White